Allison:

12(3)^3x+(-6) rewritten as 12(k)^x-2. What is the value of k?

Allison:

@Hero

Allison:

Is it 9

Hero:

Actually, I take that back.

Allison:

What

Hero:

Think more carefully about the expression you calculated to get the value of k.

Allison:

21

Allison:

Wait 9 is right

Hero:

Explain how 9 is correct.

Allison:

So it is correct

Hero:

You have not yet explained how it is correct.

Allison:

SO IT IS CORRECT

Hero:

It is not

Allison:

So it is 21

Hero:

How did you get 21? Mind explaining?

Allison:

Okay I'm just guessing -_-

Hero:

What was the expression you got for k?

Allison:

I got nothing, I'm just guessing

Allison:

Walk me through again (:

Hero:

Did you follow the process I showed you in the previous problem?

Allison:

Nope

Hero:

Well that explains it.

Allison:

:)

Allison:

>:(

Hero:

\(3^3 \ne 9\)

darkknight:

12(3)^3x+(-6) = 12(k)^x-2. We can simplify by dividing both sides by 12. 3^(3x-6) = k^(x-2) we can factor (3x-6) into 3(x-2) So basically 3^(3(x-2)) = k^(x-2) Since 3^3 is 27 then 27^(x-2) = k^(x-2) You following me so far?

darkknight:

Lmao, thanks hero. :)

Allison:

I was right the second time, it is 27

darkknight:

Actually you never said that, also it is more important knowing the process than finding the answer. Cuz if you know the process than you can always solve these types of problems

Allison:

Oh I put 27

Allison:

There is no 27 in the choices lol I'm so beat

darkknight:

Well lets talk about process, We want to simplify when we can, which is why we divided both sides by 12.

darkknight:

Now since the power on the left side is 3x-6 and the right side is x-2, we should write the left hand side so that it is also x-2, which is why we factored 3 out. Now we have 3^(3(x-2) = k^(x-2) We can rewrite this to 27^(x-2) = k^(x-2) Since they are both to the power of (x-2) then k just has to equal 27.

Allison:

Okay I understand

Allison:

Thank you for the explanation

darkknight:

np